# DCAD-2000: A Multilingual Dataset across 2000+ Languages with Data Cleaning as Anomaly Detection

### v2

@@ 6 Conclusion
In this paper, we introduce DCAD-2000, a large-scale multilingual dataset designed to address the increasing demand for high-quality and diverse training data for multilingual LLMs. Our dataset spans 2,282 languages, providing comprehensive coverage across various geographic regions, scripts (159 scripts), and larger coverage of high/medium resource languages (155 languages). To avoid manually setting thresholds during the data cleaning process, we propose a novel framework that reframes data cleaning as an anomaly detection task. This dynamic approach ensures effective identification and removal of anomalous data from noisy datasets. We evaluate DCAD-2000 on the FineTask benchmark and demonstrate the effectiveness of our proposed data cleaning framework and the high quality of the DCAD-2000 dataset.
Limitations
This work has the following limitations: (i) Although the proportion of high/medium/low resource languages in DCAD-2000 has greatly increased compare to existing multilingual datasets, a significant portion of the languages are still very low resource languages. Future work will explore to collect data for extremely low-resource languages through other modalities (e.g., images) through technologies like OCR. (ii) Due to computational limitations, we evaluate our framework and dataset only on the LLaMA-3.2-1B model. However, we expect similar evaluation results on larger models, such as 7B, 13B and even 70B models.
Ethics Statement
Our dataset integrates existing multilingual datasets, such as MaLA Ji et al. (2024) and Fineweb-2 Penedo et al. (2024b), and includes newly extracted data from Common Crawl, providing large-scale and high-quality training corpora to support the training of multilingual large language models (LLMs). Additionally, we propose a novel data cleaning method to filter out potentially toxic documents, reducing potential ethical concerns. However, performing fine-grained analysis on such a vast dataset (46.72TB) remains a significant challenge. To address this, we will release the dataset for the community to explore and research extensively. Furthermore, since our dataset is derived from open-source datasets, we will adhere to the open-source policies of these datasets to promote future research in multilingual LLMs, while mitigating potential ethical risks. Therefore, we believe our dataset does not pose greater societal risks than existing multilingual datasets.
Appendix A Statistical Analysis of the Datasets
@@ Appendix D Experimental Setup
In this section, we introduce the experimental setup used to assess the effectiveness of our proposed data cleaning method and the quality of the DCAD-2000 dataset. Our evaluation is conducted on the FineTask benchmark [^16] https://huggingface.co/spaces/HuggingFaceFW/blogpost-fine-tasks, which covers tasks in nine languages: Chinese, French, Arabic, Russian, Thai, Hindi, Turkish, Swahili, and Telugu. The benchmark spans a wide range of NLP tasks including reading comprehension, common-sense reasoning, natural language understanding, and text generation.
D.1 Task Formulations and Evaluation Metrics
Evaluation tasks in FineTask are implemented using two formulations:
  • Cloze Formulation (CF): In this setup, answer choices are not provided in the prompt, and the model directly generates each option. CF is used during the task selection phase, as it yields an early and reliable training signal.
  • Multi-Choice Formulation (MCF): Here, answer options are embedded in the prompt (e.g., using A/B/C/D prefixes), and the model is tasked with selecting the correct prefix. MCF is employed for later evaluations of pretrained models, typically using a few-shot (e.g., 5-shot) approach.
FineTask provides four evaluation metrics: Accuracy, Accuracy normalized over character length, Accuracy normalized over token length, and PMI Accuracy. However, according to statistical data, none of these metrics consistently perform well across all languages. Therefore, we chose to use normalized accuracy (norm accuracy) in our evaluation process. While base accuracy performs well in multiple-choice tasks when the option variations are small, and PMI performs best in complex reasoning and knowledge tasks, overall, length-normalized metrics (such as token- or character-based metrics) are the most reliable. However, the most suitable choice still depends on the language rather than the specific task. Therefore, we recommend selecting the maximum of character accuracy (acc_char) and token accuracy (acc_token) during evaluation. For generation tasks, unless exact matches are required, we recommend using F1 score, as it is more robust to small variations in generated outputs and less sensitive to noise.
@@ D.2 Pre-training and Evaluation Protocol
We continue pre-training the LLaMA-3.2-1B model Dubey et al. (2024) on datasets spanning the nine languages included in FineTask. Model evaluations are performed at regular intervals (approximately every 1B tokens) in a 0-shot setting (i.e., without additional prompt instructions) to capture the raw model capabilities. Normalized accuracy is used consistently as the evaluation metric. We use NVIDIA A100 40GB GPU server for our experiments and all settings are the same as the Fineweb-2 Task.
@@
Lang_Name Documents Tokens Disk Size Source
@@
Lang_Name Documents Tokens Disk Size Source
@@
Lang_Name Documents Tokens Disk Size Source
@@
Lang_Name Documents Tokens Disk Size Source

### v3

@@ 6 Conclusion
In this paper, we introduce DCAD-2000, a large-scale multilingual dataset designed to address the increasing demand for high-quality and diverse training data for multilingual LLMs. Our dataset spans 2,282 languages, providing comprehensive coverage across various geographic regions, scripts (159 scripts), and larger coverage of high/medium resource languages (155 languages). To avoid manually setting thresholds during the data cleaning process, we propose a novel framework that reframes data cleaning as an anomaly detection task. This dynamic approach ensures effective identification and removal of anomalous data from noisy datasets. Empirical experiments demonstrate the effectiveness of our proposed data cleaning framework and the high quality of the DCAD-2000 dataset across multiple multilingual benchmarks.
Appendix A Statistical Analysis of Multilingual Datasets
@@
(a) Different Languages
(b) Different Sources
(c) Different Shards
Figure 5: Distribution of average word counts across different languages, sources, and shards in the New CC dataset.
@@ Appendix D Experimental Setup
In this section, we provide a comprehensive description of our experimental setup, including dataset preparation, model configurations, continued pretraining procedure, data cleaning and anomaly detection pipeline, evaluation metrics, and implementation details.
D.1 Evaluation Benchmarks
FineTask Benchmark.
FineTask is a multilingual, multi‑task benchmark covering nine typologically diverse languages: Chinese, French, Arabic, Russian, Thai, Hindi, Turkish, Swahili, and Telugu. The benchmark spans a wide range of NLP tasks including reading comprehension, common-sense reasoning, natural language understanding, and text generation. FineTask provides four evaluation metrics: Accuracy, Accuracy normalized over character length, Accuracy normalized over token length, and PMI Accuracy. However, according to statistical data, none of these metrics consistently perform well across all languages. Therefore, we chose to use normalized accuracy (norm accuracy) in our evaluation process.
Multilingual Benchmarks.
To analyze performance across varying resource levels, we further evaluate on three established multilingual corpora:
  • SIB‑200 [24]: A suite of topic classification datasets across 205 languages. We use the raw accuracy on held‑out test sets.
  • Glot500‑c [8]: A curated corpus spanning 500 languages for generation and language modeling. We compute negative log‑likelihood (NLL) on held‑out sentences:
    $\mathrm{NLL}=-\frac{1}{T}\sum_{t=1}^{T}\log p_{\theta}(w_{t}\mid w_{<t}),$ (9)
    where $T$ is the total token count in the evaluation set.
  • FLORES‑200 [25]: A benchmark for low‑resource machine translation covering 200 languages. We translate from English into each target language (Eng-XX) and translate from other languages into English (X-Eng) and evaluate using SacreBLEU with default settings.
@@ D.2 Pre-training and Evaluation Protocol
We perform continued pretraining on three representative decoder-only large language models (LLMs): LLaMA‑3.2‑1B [47], Qwen‑2.5‑7B [48], and Aya‑expanse‑32B [49]. These models are selected to represent a diverse range of open-source models across different parameter scales, allowing us to investigate the effects of the different dataset, data cleaning pipeline across small, medium, and large model sizes. All models are accessed and managed through the HuggingFace Transformers library.
Given the limitations of computational resources, we refrain from full-parameter finetuning. Instead, we adopt Low-Rank Adaptation (LoRA; 53), a parameter-efficient fine-tuning technique that introduces trainable low-rank matrices into each transformer layer, substantially reducing the number of trainable parameters while maintaining competitive performance.
Our training pipeline closely follows the setup described in the LightEval repository [^16] https://github.com/huggingface/lighteval, a lightweight evaluation and fine-tuning framework developed by HuggingFace. This ensures reproducibility and consistency with widely adopted community practices. Experiments are conducted on NVIDIA A100 GPU with 80GB memory, which provides sufficient memory bandwidth and compute capability to support batch-level parallelism and efficient LoRA-based fine-tuning. All hyperparameters and task-specific configurations are aligned with those used in the Fineweb-2 benchmark, ensuring comparability with previous work and consistent evaluation conditions.
D.3 Statistical Anomaly Detection
We provide a detailed comparison of the anomaly detection algorithms evaluated for data cleaning in DCAD-2000. The methods are selected based on their popularity, conceptual diversity, and availability in scikit-learn [^17] https://scikit-learn.org. All experiments are conducted using the same eight-dimensional feature vectors described in Section 3.
Isolation Forest (ISO_Forest)
[46] is an ensemble-based method that isolates anomalies instead of profiling normal data points. It constructs random binary trees by recursively selecting features and split values, and then uses the path length of each data point across the trees to assess anomaly scores. Shorter paths indicate higher likelihood of being an outlier. ISO_Forest is well-suited to high-dimensional and noisy data and requires minimal hyperparameter tuning. Its main drawback is higher computational cost relative to simpler methods, though it scales well with the number of samples.
One-Class SVM (OC_SVM)
[50] is a kernel-based method that attempts to separate the data from the origin in a transformed feature space. It is sensitive to the choice of kernel function (e.g., RBF, linear) and associated parameters (e.g., gamma, nu). OC_SVM can be effective in capturing complex boundaries, but it often suffers from scalability issues and requires careful parameter tuning, especially in high-dimensional multilingual settings like DCAD-2000.
Local Outlier Factor (LOF)
[51] is a density-based method that identifies anomalies based on local density deviation. It compares the local density of a data point with that of its neighbors. Points that have substantially lower density than their neighbors are considered outliers. The performance of LOF depends heavily on the number of neighbors chosen and tends to degrade in high-dimensional spaces due to the curse of dimensionality. It is also computationally expensive for large datasets.
K-Means
[52] is a clustering algorithm typically used for unsupervised partitioning of data. For anomaly detection, it is repurposed by measuring the distance of points from their assigned cluster centroids—points that are far from any centroid can be considered anomalous. K-Means is computationally efficient and easy to implement but lacks sensitivity to local structures and does not inherently model outliers. Its effectiveness depends on a suitable choice of the number of clusters.
Appendix E Limitations
This work has the following limitations: (i) Although the proportion of high/medium/low resource languages in DCAD-2000 has greatly increased compare to existing multilingual datasets, a significant portion of the languages are still very low resource languages. Future work will explore to collect data for extremely low-resource languages through other modalities (e.g., images) through technologies like OCR. (ii) We evaluate the new data cleaning framework only on four classical anomaly detection algorithms; however, since the framework is algorithm-independent, it should also be effective with other anomaly detection algorithms. (iii) For language identification, we use GlotLID [21], a FastText-based model whose limitations in handling massive multilinguality have been discussed in previous works [54]. However, since the data cleaning pipeline is language-agnostic, other language identification models can also be employed.
Appendix F Ethics Statement
Our dataset integrates existing multilingual datasets, such as MaLA [11] and Fineweb-2 [14], and includes newly extracted data from Common Crawl, providing large-scale and high-quality training corpora to support the training of multilingual large language models (LLMs). Additionally, we propose a novel data cleaning method to filter out potentially toxic documents, reducing potential ethical concerns. However, performing fine-grained analysis on such a vast dataset (46.72TB) remains a significant challenge. To address this, we released the dataset for the community to explore and research extensively. Furthermore, since our dataset is derived from open-source datasets, we adhere to the open-source policies of these datasets to promote future research in multilingual LLMs, while mitigating potential ethical risks. Therefore, we believe our dataset does not pose greater societal risks than existing multilingual datasets.
@@ Appendix F Ethics Statement
Lang Code Documents Tokens Disk Size Source
@@
Lang Code Documents Tokens Disk Size Source
@@
Lang Code Documents Tokens Disk Size Source
@@
Lang Code Documents Tokens Disk Size Source